0.2 kg of water at 59oC is added to 0.4 kg of water at 20oC in a polystyrene cup. What is
the final temperature of the mixture?
(s.h.c. of water is 4200 J/kg/oC )
1
Expert's answer
2020-11-14T14:00:27-0500

Solution.

"Q1(H2O) = Q2(H2O)"

"c \times m1 \times (T1-T2) = c \times m2 \times (T2-T1)"

"0.2 \times (59-x) = 0.4 \times (x-20)"

x = 33

Answer:

T2 = 33 Celsius degree